You are a graduating mechanical engineering student. As part of the course requirement, you applied for an internship in a model makers company. The officer – in – charge asked your assistance to make a model of a car for a client. The dimension of the car is 5 ft by 10 ft. Based on this, you are required to make a model with a scale of 1 in : 4 ft.


1
Expert's answer
2021-07-01T06:53:23-0400

A model is the particular vehicle type that an automaker produces.

Dimension of the car

5ft by 10ft

model scale

1 in: 4ft


"4ft=1inch\\

5ft=\frac{5\times 1}{4}\\=1.25 \space inch"


"10ft=\frac{10}{4}\\=2.5\space inch"


Dimension the constructed model is

1.25 inch by 2.5 inch